OpenStudy (anonymous):

easy question- 10 % × 55 = _____

Parth (parthkohli):

5.5

OpenStudy (anonymous):

10%=10/100=0.1 0.1x55=5.5

Parth (parthkohli):

1/10 * 55 = 55/10 = 5.5

OpenStudy (anonymous):

move the decimal 2 times to the left than times it by 55

OpenStudy (anonymous):

ii wrote the wrong problem my bad